Grade 12Magical Mathematics[Interesting Approach]

If Ar, r°1,2,3,.....,n be points on a circle of unit radius then prove that the sum of thair mutual distance is less than n².

Arindam Santra
7 Years agoGrade 12

1 Answer

rahul
7 Years ago
Imagine it this way, the maximum distance between any two points in the circle is equal to its diameter.So if n=2 then the maximum distance between the points is 2, which is less than n= (4).
Now imagine that there is an n-sided polygon inscribed in a circle of unit radius,then the maximum sum of all distances between them would be nC*2 =n(n-1)=n2-12, for all n,